# Modeling information spread across networks with communities using a multitype branching process framework

### v1

@@ Appendix B Recovering probabilities from probability generating functions
In this section we give detailed explanation on how the probabilities are recovered from the corresponding probability generating functions. As was explained in Sec. III, the probability generating function, $G_{\boldsymbol{N}(t)}$, by definition is
@@ Appendix C Comparing branching process simulations to network simulations
We compare our model using multi-type branching processes to network-based simulations and to the expected information diffusion using the mean-matrix to check the accuracy of the model. The mean-matrix $M$ tracks the expected offspring of each type, so $M_{ij}$ is the expected number of offspring from a node $i$ that are of type $j$ in the next time step. For the SBM we have the mean-matrix
@@
Figure 8 shows the results obtained for $10,000$ simulations for a range of values of infection probability $\rho$ ($\rho=0.04$, $\rho=0.06$, $\rho=0.08$), expected number of edges inside a community $\lambda_{\text{in}}$ ($\lambda_{\text{in}}=4$, $\lambda_{\text{in}}=8$) and expected number of edges between communities $\lambda_{\text{out}}$ ($\lambda_{\text{out}}=0$, $\lambda_{\text{out}}=2$, $\lambda_{\text{out}}=4$) on a Poisson-distributed network. We use Stochastic Block Model networks containing $10,000$ nodes split evenly between community 1 community 2 and compare those to the equivalent multi-type branching process model for Poisson distribution. The simulations were set to start with the activation of one random node in community 1. There is cross-community information spread when $\lambda_{\text{out}}>0$, in which case community 2 is expected to be infected on the first passage, on average. Fig. 8c) shows the offspring distribution for $\rho=0.08$, $\lambda_{\text{in}}=8$ and $\lambda_{\text{out}}=2$. This is the combination of parameters among the considered values that generates the longest process. The process dies off earlier for the simulations using $\rho=0.06$ (Fig. 8b)) and $\rho=0.04$ (Fig. 8a)) since the information survives shortly when the probability of infection is lower. When we consider no possibility of cross-community information spread ($\lambda_{\text{out}}=0$, Fig. 8f)), we observe diffusion only inside community 1, and the overall process dies off earlier than the analogous with $\lambda_{\text{out}}=2$ (Fig. 8b)). On the other hand, when we increase $\lambda_{\text{out}}$ (Fig. 8d)) information spreads further in both community 1 and community 2 in comparison to the analogous with lower $\lambda_{\text{out}}$ values (Fig. 8b) and f)). When we keep $\lambda_{\text{out}}$ moderate ($\lambda_{\text{out}}=4$) and decrease $\lambda_{\text{in}}$ ($\lambda_{\text{in}}=\lambda_{\text{out}}=4$), with moderate probability of infection, the information dies out early.
@@ Appendix D Log-normal network parameter sweep
In this section we provide a parameter sweep for both $\rho$ and $p_{e}$, and analyse the resultant distributions of cascade sizes. The network is the same as that presented in Sec. IV, namely where we have two communities of $5000$ nodes, each with a log-normal distribution internally. Community 1 has $\mu_{1}=0.85$ and $\sigma_{1}=0.4$, resulting in an internal average degree of 3. Community 2 has $\mu_{2}=1.8$ and $\sigma_{2}=0.4$, resulting in an internal average degree of 7. The cross-community links are selected by adding single edges between nodes. We randomly select one node from community 1 and one node from community 2 and link them. The proportion of nodes selected is $p_{e}$.
@@
Figure 9 shows the comparison of the theory and simulation under different values of $p_{e}$, where $\rho$ is fixed. We can see that we have excellent agreement between our theory and simulations. As we increase the number of cross-community edges we have a general increase in cascade size. This is most pronounced for those cascades seeded in community 1 (solid line).
@@
Figure 9: Parameter sweep, where $\rho$ is fixed at $0.1$ and $p_{e}$ is swept from $10/5000$ to $2500/5000$. Each community has 5000 nodes, totalling 10000 nodes in the network. Lines represent theory curves, and the points represent the results of $N_{sim}=10^{6}$ Monte Carlo simulations of the ICM model on the network. Note that the middle dot-dashed line is for the corresponding simple branching process (which does not take into account community structure).
@@
Figure 10 provides the behaviour of our cross-community branching processes for various $\rho$ values, where $p_{e}$ is fixed to $1000/5000$. We can note that we have excellent agreement between our theory and simulations in general. For larger $\rho$ (most pronounced for $\rho=0.14$) values, we can note discrepancies between the theory and simulated values. The larger simulated cascade do not reach the same sizes as the theoretical cascades sizes. This is due to the finite size of the network, as for this value, cascades activate a significant proportion of the networks, while in the theory it is assumed that the networks are infinitely large.
@@
Figure 10: Parameter sweep, where $p_{e}$ is fixed at $1000/5000$ and $\rho$ is swept from $0.005$ to $0.14$. Each community has 5000 nodes, totalling 10000 nodes in the network. Lines represent theory curves, and the points represent the results of $N_{sim}=10^{6}$ Monte Carlo simulations of the ICM model on the network. Note that the middle dot-dashed line is for the corresponding simple branching process (which does not take into account community structure).
@@ Appendix E Constructing a multi-type branching process for k-regular networks
from which we can easily show that the excess offspring distributions for between-community spread are given by $G_{\widetilde{X}_{2}^{(1)}}(s)=G_{\widetilde{X}_{1}^{(2)}}(s)=1$. Now we can proceed to create the offspring distribution for each of the 4 types that we have to track in our branching process. First, let us consider the PGF for the offspring from a node in community 1, who was activated from within its community. As it was activated from inside community 1, it will have one less neighbour that it can activate and this requires the use of the excess distribution, which is associated random variable $\widetilde{X}_{1}^{(1)}$, while no cross-community links we used in the activation, and so will use the cross-community degree distribution, which is associated with the random variable $X_{2}^{(1)}$, yielding the following offspring distribution:
@@
To find the number of nodes active at time $t$ we just need to consider what happens between generation 0 and 1. If we seed the process with one active individual in community 1 at time $t=0$, then our initial conditions are $G_{\mathcal{N}(1)}(s)=(s)^{1}$. Then as the process has just started the active node has no previously active neighbours, therefore, the offspring distribution is given by
@@
and this process is repeated until the required generation has been reached. We can then marginalised the required types in the PGF to recover the univariate PGF for the cascade size distribution, for example, by $G_{\mathcal{C}(t)}(1,1,1,1,c)$. Then this can be inverted via inverse fast Fourier transform to obtain the probability mass function, see App. 39. It should be noted that this entire process is no more complex for general degree distributions once we have the network structure, just a little more cumbersome.

### v2

@@ Appendix B Recovering probabilities from probability generating functions
In this section we give detailed explanation on how the probabilities are recovered from the corresponding probability generating functions. As explained in Sec. III, the probability generating function, $G_{\boldsymbol{N}(t)}$, by definition is
@@ Appendix C Comparing branching process simulations to network simulations
We compare our model using multi-type branching processes to network-based simulations and to the expected diffusion using the mean-matrix to check the accuracy of the model. The mean-matrix $M$ tracks the expected offspring of each type, so $M_{ij}$ is the expected number of offspring from a node $i$ that are of type $j$ in the next time step. For the SBM we have the mean-matrix
@@
Figure 8 shows the results obtained for 10 00010\,000 simulations for a range of values of infection probability $\rho$ ($\rho=0.04$, $\rho=0.06$, $\rho=0.08$), expected number of edges inside a community $\lambda_{\text{in}}$ ($\lambda_{\text{in}}=4$, $\lambda_{\text{in}}=8$) and expected number of edges between communities $\lambda_{\text{out}}$ ($\lambda_{\text{out}}=0$, $\lambda_{\text{out}}=2$, $\lambda_{\text{out}}=4$) on a Poisson-distributed network. We use Stochastic Block Model networks containing 10 00010\,000 nodes split evenly between community 1 community 2 and compare those to the equivalent multi-type branching process model for Poisson distribution. The simulations were set to start with the activation of one random node in community 1. There is cross-community spread when $\lambda_{\text{out}}>0$, in which case community 2 is expected to be infected on the first passage, on average. Fig. 8c) shows the offspring distribution for $\rho=0.08$, $\lambda_{\text{in}}=8$ and $\lambda_{\text{out}}=2$. This is the combination of parameters among the considered values that generates the longest process. The process dies off earlier for the simulations using $\rho=0.06$ (Fig. 8b)) and $\rho=0.04$ (Fig. 8a)) since the infection survives shortly when the probability of infection is lower. When we consider no possibility of cross-community diffusion ($\lambda_{\text{out}}=0$, Fig. 8f)), we observe diffusion only inside community 1, and the overall process dies off earlier than the analogous with $\lambda_{\text{out}}=2$ (Fig. 8b)). On the other hand, when we increase $\lambda_{\text{out}}$ (Fig. 8d)) infection spreads further in both community 1 and community 2 in comparison to the analogous with lower $\lambda_{\text{out}}$ values (Fig. 8b) and f)). When we keep $\lambda_{\text{out}}$ moderate ($\lambda_{\text{out}}=4$) and decrease $\lambda_{\text{in}}$ ($\lambda_{\text{in}}=\lambda_{\text{out}}=4$), with moderate probability of infection, the process dies out early.
@@ Appendix D Log-normal network parameter sweep
In this section we provide a parameter sweep for both $\rho$ and $p_{e}$, and analyze the resultant distributions of cascade sizes. The network is the same as that presented in Sec. IV, namely where we have two communities of 5 0005\,000 nodes, each with a log-normal distribution internally. Community 1 has $\mu_{1}=0.85$ and $\sigma_{1}=0.4$, resulting in an internal average degree of 3. Community 2 has $\mu_{2}=1.8$ and $\sigma_{2}=0.4$, resulting in an internal average degree of 7. The cross-community links are selected by adding single edges between nodes. We randomly select one node from community 1 and one node from community 2 and link them. The proportion of nodes selected is $p_{e}$.
@@
Figure 9 shows the comparison of the theory and simulation under different values of $p_{e}$, where $\rho$ is fixed. We can see that we have excellent agreement between our theory and simulations. As we increase the number of cross-community edges we have a general increase in cascade size. This is most pronounced for those cascades seeded in community 1 (purple line).
@@
Figure 9: Parameter sweep, where $\rho$ is fixed at $0.1$ and $p_{e}$ is swept from 10/5 00010/5\,000 to 2 500/5 0002\,500/5\,000. Each community has 5 0005\,000 nodes, totaling 10 00010\,000 nodes in the network. Lines represent theory curves, and the points represent the results of $N_{sim}=10^{6}$ Monte Carlo simulations of the ICM model on the network. Note that the middle pink line is for the corresponding simple branching process (which does not take into account community structure).
@@
Figure 10 provides the behavior of our cross-community branching processes for various $\rho$ values, where $p_{e}$ is fixed to 1 000/5 0001\,000/5\,000. We can note that we have excellent agreement between our theory and simulations in general. For larger $\rho$ (most pronounced for $\rho=0.14$) values, we can note discrepancies between the theory and simulated values. The larger simulated cascade do not reach the same sizes as the theoretical cascades sizes. This is due to the finite size of the network, as for this value, cascades activate a significant proportion of the networks, while in the theory it is assumed that the networks are infinitely large.
@@
Figure 10: Parameter sweep, where $p_{e}$ is fixed at 1 000/5 0001\,000/5\,000 and $\rho$ is swept from $0.005$ to $0.14$. Each community has 5 0005\,000 nodes, totaling 10 00010\,000 nodes in the network. Lines represent theory curves, and the points represent the results of $N_{sim}=10^{6}$ Monte Carlo simulations of the ICM model on the network. Note that the middle pink line is for the corresponding simple branching process (which does not take into account community structure).
As shown in Fig. 10, cascades seeded in the second community appear to follow a power-law distribution. Table 2 below illustrates the power-law exponent ($\alpha$) derived using a maximum likelihood approach with the simulated cascade data for multiple infection probabilities ($\rho$). Notably, as the infection probability increases and the process nears the critical point, the characteristic exponent approaches $3/2$, which is consistent with expectations for the critical regime.
Estimating the power-law exponent for cascades seeded in the first community is more complex due to a sharp transition between two regimes in their cascade size distribution. The cascades modeled using the SBP in the figure primarily serve to demonstrate that the SBP model is not well-suited for diffusion processes on networks with communities. As such, these results were not pursued further. It is important to note that these results are derived from the simulation data. Ideally, we would like to derive asymptotic expression for the cascade size distribution to predict this scaling; this is left to further work.
Cascades seeded in community 2
$\rho$ 0.08 0.10 0.12 0.14
$\alpha$ 1.83 1.72 1.60 1.54
Table 2: Estimates of the scaling parameter $\alpha$ for different infection probabilities $\rho$ obtained using the maximum likelihood estimator method newman_networks_2018. All other simulation parameters are the same as those in Fig. 10.
@@ Appendix E Constructing a multi-type branching process for k-regular networks
from which we can easily show that the excess offspring distributions for between-community spread are given by $G_{\widetilde{X}_{2}^{(1)}}(s)=G_{\widetilde{X}_{1}^{(2)}}(s)=1$. Now we can proceed to create the offspring distribution for each of the 4 types that we have to track in our branching process. First, let us consider the PGF for the offspring from a node in community 1, who was activated from within its community. As it was activated from inside community 1, it will have one less neighbor that it can activate and this requires the use of the excess distribution, which is associated random variable $\widetilde{X}_{1}^{(1)}$, while no cross-community links we used in the activation, and so will use the cross-community degree distribution, which is associated with the random variable $X_{2}^{(1)}$, yielding the following offspring distribution:
@@
To find the number of nodes active at time $t$ we just need to consider what happens between generation 0 and 1. If we seed the process with one active individual in community 1 at time $t=0$, then our initial conditions are $G_{\mathcal{N}(1)}(s)=(s)^{1}$. Then as the process has just started the active node has no previously active neighbors, therefore, the offspring distribution is given by
@@
and this process is repeated until the required generation has been reached. We can then marginalized the required types in the PGF to recover the univariate PGF for the cascade size distribution, for example, by $G_{\mathcal{C}(t)}(1,1,1,1,c)$. Then this can be inverted via inverse fast Fourier transform to obtain the probability mass function, see App. 39. It should be noted that this entire process is no more complex for general degree distributions once we have the network structure, just a little more cumbersome.
Appendix F The community seeding effect on a log-normal network with communities of similar degree distribution
This section repeats the experiments from Sec. IV.2, where we isolated the effect of inter community spreading by comparing two communities with approximately log normally degree distribution, where community 1 had an average degree of approximately 3 and community 2 had an approximate degree of 7. We were able to note in Fig. 7 that community 1, with the lower average degree, had a qualitatively different cascade distribution when we allowed inter community spread; whereas community 2 had no such discernible effect. Here we repeat this comparison but now using two communities with closer average degrees; now community 1 has an average degree of 4 and community 2 has an average degree of 5.
We can note from Fig. 11, that again we have excellent agreement between our simulation and our theory curves. In Fig. 11 it is hard this time to see any qualitative change in the shape of the cascade size distributions in community 1 or community 2 like we saw before.
However, we can still isolate any effect that we might see by comparing the cascade size distribution in community 1 when we allow inter community spreading ($p_{e}=0.02$) and when we do not ($p_{e}=0$) in Fig. 12a) and 12b). Again, it is hard to discern a difference with the simulations masking any patterns; However, if we focus on the theory curves with no simulations and examine a wider range of probabilities, shown in Fig. 12b), we can see a similar difference emerging. This difference occurs at a significantly smaller probability than previously, which means that even when community 1 does see a benefit from spreading to community 2, which has a slightly higher average degree, it only becomes pronounced at a very small probability, i.e., it would be rare to observe this effect for this network configuration. Again, we can isolate any effect that we might see by comparing the cascade size distribution in community 2 when we allow inter community spreading ($p_{e}=0.02$) and when we do not ($p_{e}=0$) in Fig. 12c). Like before, we cannot see a difference; which means that there is little qualitative difference when we allow spreading between communities in the case where we seed cascades in community 2.
Figure 11: Cascade size across the entire network for a range of infection ($\rho$) parameters when seeded in different communities, leaving the cross-community connection probability, $p_{e}$, fixed at $0.02$. Each community has 5 0005\,000 nodes, totaling 10 00010\,000 nodes in the network. Community 1 has an internal average degree of $4$, while community 2 has an internal average degree of $5$. Note that the middle pink line is for the corresponding simple branching process (which does not take into account community structure).
Figure 12: Cascade size across the entire network where $\rho$ is fixed to $0.16$. Purple and blue lines (MTBP) represent $p_{e}=0.02$, where cross-community spread is possible, and pink lines (SBP) represent $p_{e}=0$, where spread is confined to one community. a) Diffusion behavior when seeding in community 1 (which has an average degree of $\approx 4$). b) Diffusion behavior when seeding in community 1 (theory curves only to show the dynamics on cascades that spread further). c) Diffusion behavior when seeding in community 2 (which has an average degree of $\approx 5$). Here the blue and pink lines coincide. As before, each community has 5 0005\,000 nodes, totaling 10 00010\,000 nodes in the network.